OpenStudy (anonymous):

ᶴ(cos x/3sinx+4√(sin x)) please i need procedure how to integrate this function

OpenStudy (dumbcow):

lets see , try u sub then partial fractions u = sin x du = cos x dx that will get rid of the cos x on top then \[\int\limits \frac{du}{3u + 4\sqrt{u}} = \int\limits \frac{du}{\sqrt{u} (3 \sqrt{u} +4)}\] \[\rightarrow \frac{A}{\sqrt{u}} + \frac{B}{3\sqrt{u} +4} = \frac{1}{\sqrt{u}(3 \sqrt{u} +4)}\] 3A +B = 0 4A = 1 --> A = 1/4 --> B = -3/4 \[\frac{1}{4} \int\limits \frac{du}{\sqrt{u}} - \frac{3}{4} \int\limits \frac{du}{3 \sqrt{u} + 4}\]